# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v2

@@ 4.1 Discrete Diffusion Divergence Instruction
A primary challenge in distilling MDMs is the discrete nature of the state space. The gradient formulation in Luo et al. (2023b) relies on differentiating through $\mathbf{z}_{t}$. In MDMs, this forward process involves non-differentiable operations (e.g., $\operatorname*{arg\,max}$), which renders this gradient inapplicable (Zekri and Boullé, 2025). Zhu et al. (2025b) proposed a proxy model to approximate $\mathbf{z}_{t}$ and achieved promising results in text-to-image tasks. However, its effectiveness in text generation remains underexplored. We instead take inspiration from the policy gradient (Schulman et al., 2017; Fan et al., 2023) to adapt a mathematically rigorous solution. By decomposing the gradient of the objective as a score function weighted by a log-density ratio, we derive the following gradient estimation.
@@
Let the objective $\mathcal{L}(\nu)$ be the weighted integral of KL divergences between student and teacher marginals, $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. The gradient of the objective admits:
@@
where the expectation over time $t$ is sampled from distribution $\pi(t)$, and $R(\mathbf{z}_{t},t):=\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$ denotes the reward (log-density ratio between student and teacher) at $\mathbf{z}_{t}$.
The proof is deferred to Appendix C. Theorem 4.1 shows that the IKL gradient admits a score-function form that does not differentiate through the discrete sampling path. We sample $\mathbf{x}\sim\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)$, $\mathbf{z}_{t}\sim\mathcal{Q}$ and differentiate only $\log\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)$. The reward $R(\mathbf{z}_{t},t)$ then weights this score to guide distillation toward matching the teacher’s marginal at time $t$.
@@ 4.2 Density‑Ratio Estimation
The reward is a mathematically natural choice for distribution matching. However, both $\log\mathbf{q}_{\nu}$ and $\log\mathbf{q}_{\theta}$ are intractable to compute directly. To overcome this, we avoid modeling the individual densities and instead approximate their ratio. Inspired by Wang et al. (2025) in continuous-time diffusion, we train an auxiliary discriminator specifically for this it. See Appendix D for details.
@@
Let $D_{\lambda}:\mathcal{V}^{L}\times[0,1]\to(0,1)^{L}$ be a parameterized discriminator to distinguish samples generated by $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. For the optimal discriminator $D_{\lambda^{\star}}(\mathbf{z}_{t},t)$, the density ratio can be expressed as
@@
Following Lemma 4.2, we subsequently construct a tractable reward signal $R(\mathbf{z}_{t},t)$ by aggregating the log-density ratios over all masked positions. With $M$ denoting the number of masked tokens in the sequence, the reward can be estimated from the discriminator network $D_{\lambda}$ as follows:
@@
where $D_{\lambda}^{\ell}(\cdot,\cdot)$ denotes the $\ell$-th element of $D_{\lambda}(\cdot,\cdot)$ ($\ell=1,2,\cdots,L$). This procedure provides a tractable and computable reward signal for our objective according to $D_{\lambda}$.
@@ 4.3 Grouped Reward Normalization
While the reward estimator in (6) is tractable, its direct use in score-function gradients can exhibit high variance (Williams, 1992). We therefore adopt Group Relative Policy Optimization (Shao et al., 2024; Team et al., 2025) to standardize rewards within each mini-batch. Given a mini-batch of size $G$, we draw $\{t_{i}\}_{i=1}^{G}$ from $\pi(t)$, estimate $\{\mathbf{z}_{t_{i}}\}_{i=1}^{G}$ and $\{R(\mathbf{z}_{t_{i}},t_{i})\}_{i=1}^{G}$ correspondingly [^3] For clarity and concision, we denote $R(\mathbf{z}_{t_{i}}^{1:L},t_{i})$ as $R_{i}$, and $D_{\lambda}(\mathbf{z}_{t_{i}}^{1:L},t_{i})$ as $D_{i}$ throughout.. The final stabilized reward $\widetilde{R}_{i}$ is obtained by normalizing $R_{i}$ with the mean $\mu_{g}$ and variance $\sigma_{g}^{2}$:
$\widetilde{R}_{i}=\frac{R_{i}-\mu_{g}}{\sigma_{g}+\epsilon},\qquad\text{ where }\mu_{g}=\tfrac{1}{G}\sum_{i=1}^{G}R_{i},\text{ and }\sigma_{g}^{2}=\tfrac{1}{G}\sum_{i=1}^{G}(R_{i}-\mu_{g})^{2}.$ (7)
Here, a small constant $\epsilon>0$ is used for numerical stability. By substituting this stabilized reward $\tilde{R}_{i}$ for the raw reward $R_{i}$ in our gradient estimator, we arrive at a more robust update rule.
@@ 4.4 The Proposed Algorithms
Adversarial reward estimation. Following Wang et al. (2025), we train a parameterized discriminator $D_{\lambda}$ to separate corrupted samples from $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ at time $t$. For each mini-batch index $i$, we start from the all masked sequence to sample $\mathbf{x}\sim\mathbf{p}_{\nu}$ and $\mathbf{x}^{\prime}\sim\mathbf{p}_{\theta}$. We then corrupt both to the same noise level $t_{i}$ via the forward process (2) to obtain $\mathbf{z}_{i}\sim\mathcal{Q}$ and $\mathbf{z}^{\prime}_{i}\sim\,\mathcal{Q}$. The discriminator is optimized with balanced binary cross-entropy:
@@
At optimality $\lambda^{\star}$, $D_{\lambda^{\star}}(\mathbf{z}_{t},t)\approx\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t},t)}{\mathbf{q}_{\nu}(\mathbf{z}_{t},t)+\mathbf{q}_{\theta}(\mathbf{z}_{t},t)}$, which indicates that $\mathrm{logit}\,D_{\lambda^{\star}}$ can provide a tractable estimate of the log–density ratio used in (6).
@@
Figure 2: The pipeline of DiDi-Instruct (Algorithm 1). Given a fully masked input $\mathbf{z}_{t}$ ($t=1$), both the student $\mathbf{p}_{\nu}$ and the teacher $\mathbf{p}_{\theta}$ produce clean samples $\mathbf{x}$ and $\mathbf{x}^{\prime}$, which are corrupted at $t_{i}\sim\pi(t)$ to form $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{\prime}$. The discriminator $D_{\lambda}$ is trained to classify these outputs, while its reward signal (6) enables the gradient update (5) for the student. The red line denotes the gradient flow for the student’s update step, and the blue line represents the one for the auxiliary model’s update step.
End-to-end training algorithm. The DiDi-Instruct training procedure alternates between updating the discriminator $D_{\lambda}$ and the student $\mathbf{p}_{\nu}$, which is summarized in Algorithm 1 and visualized in Figure 2. Each iteration consists of two phases: first, the discriminator is updated to better distinguish between corrupted samples between $\mathbf{p}_{\theta}$ and $\mathbf{p}_{\nu}$; we apply (5) with the normalized reward and the decomposed score to stabilize gradient updates. This alternating scheme yields a robust few-step generator that closely matches the teacher’s marginals over different corrupt levels.
Algorithm 1 DiDi-Instruct
for each training step do
  Sample $t_{i}\sim\pi(t)$, and
   $\mathbf{x}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{\nu}}(\mathbf{z}_{t},t=1)$, $\mathbf{x}^{\prime}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{\theta}}(\mathbf{z}_{t},t=1)$.
  Corrupt $\mathbf{x}$, $\mathbf{x}^{\prime}$ to partially masked $\mathbf{z}_{i}$, $\mathbf{z}^{\prime}_{i}$.
  Update discriminator $D_{\lambda}$ with (8).
  Update student $\mathbf{p}_{\nu}$ with (5)-(7), and (9).
return student $\mathbf{p}_{\nu}$ and discriminator $D_{\lambda}$
Algorithm 2 RGAS
Initialize $\mathbf{z}_{N}\leftarrow(\mathbf{m},\dots,\mathbf{m})$.
for $n=N,\dots,1$ do
  if $n$ in early stage then
    Sample $\mathbf{z}_{n-1}$ with (10) ($h>0$, $M=1$).
  else
    Sample $\mathbf{z}_{n-1}$ with (10)-(11) ($h=0$, $M\geq 1$).
Set $\mathbf{x}=\mathbf{z}_{0}$ and return sequence $\mathbf{x}$
@@
where the logits for unmasked tokens are augmented with the reward gradient:
@@
Here, $\mathbf{e}_{z_{n}^{\ell}}\in\mathcal{V}$ denotes the one-hot vector with a $1$ at index $z_{n}^{\ell}$ and 0 elsewhere, and $h$ denotes the tilting scale. RGAS adjusts $h$ and the number of candidates $M$ across the denoising process: For early steps ($t_{n}\approx 1$), we use gradient tilting ($h>0,M=1$) to steer global structure toward high-reward regions. For late steps ($t_{n}\approx 0$), we switch to multi-candidate re-ranking ($h=0,M>1$): we draw $M$ candidates $\{\mathbf{z}_{n-1}^{(m)}\}_{m=1}^{M}$ [^4] For simplicity, we denote the $m$-th candidate of $\mathbf{z}_{t_{n}}^{1:L}$ as $\mathbf{z}_{n}^{(m)}$. from (10) with $h=0$, then select $\mathbf{z}_{n-1}$ following:
@@
To decompose score estimation in (9) and sample the intermediate state $\mathbf{z}_{i}$ at $t_{i}$, we set $h=0$ and $M=1$ to mitigate the risk of reward hacking (Skalse et al., 2022; Gao et al., 2023). The complete procedure is detailed in Algorithm 2.
@@ 5 Experiments
Our experiments are designed to distill a pre-trained teacher model into an efficient few-step student generator with DiDi-Instruct. All models are trained on OpenWebText (Gokaslan et al., 2019). Following standard practices (Sahoo et al., 2025), we tokenize the corpus using the GPT-2 tokenizer, pack sequences to a context length of 1024, and hold out the last 100,000 documents for validation.
Our teacher is a 169M parameter MDLM with a Diffusion Transformer (Peebles and Xie, 2023) (12 layers, 12 attention heads, 768 hidden dimension). We pre-trained this model for 1024 NFEs, achieving a perplexity of 38.53 and an entropy of 5.22. The student model shares an identical architecture to ensure a fair comparison. The discriminator is a 131M parameter network based on the same backbone, but with a new randomly initialized classification head. This head consists of two linear layers with spectral normalization and a SiLU activation function. During distillation, the discriminator model and the student generator are trained adversarially in an alternating fashion.
The distillation process runs 10,000 iterations using the AdamW optimizer with a learning rate of $10^{-6}$ and no warm-up. The teacher model was pre-trained on 8 NVIDIA H100 GPUs. In contrast, our DiDi-Instruct distillation is highly efficient, requiring only a single H100 GPU. All training procedures leverage bfloat16 for acceleration. For more details, please refer to Appendix E.
@@ 5.1 Experimental Result and Analysis
We present the empirical results of our proposed DiDi-Instruct. The results show that DiDi-Instruct not only achieves state-of-the-art performance in sample quality, particularly in the few-step generation regime, but also offers substantial improvements in training and inference efficiency.
Generation results. We evaluate generative quality by sampling text from the distilled student from $8$ to $128$ NFEs and measuring GPT-2 Large generative PPL and average sequence entropy. Figure 1 shows that DiDi-Instruct consistently outperforms all baselines in PPL across all sampling steps. Notably, with only 16 NFEs, our model’s PPL already surpasses that of the 1024-step teacher model. At 1024 NFEs, DiDi-Instruct achieves a final PPL of 15.62, a reduction of over 24% compared to the strongest baseline. These performance gains are achieved with a negligible loss in diversity; the generative entropy from 8 to 128 NFEs is 5.17, 5.22, 5.20, 5.18, and 5.17, respectively (e.g., 1024-step teacher model is 5.22), indicating sample diversity is well-preserved.
Efficiency-performance tradeoff. DiDi-Instruct offers substantial computational advantages in both training and inference. Our single-round distillation framework completes training in around one H100 GPU hour, in contrast to the 2020+ GPU hours required by multi-round methods (Sahoo et al., 2024; Deschenaux and Gulcehre, 2025). During inference, RGAS provides a superior perplexity-latency trade-off. As shown in Figure 3, it achieves a PPL of $23.54$ at a latency of $1.31$ s/sequence ($32$ NFEs), whereas AS only reaches a PPL of $44.54$ ($0.71$ s/sequence, $32$ NFEs).
@@
Sample quality. Appendix F.7 presents full text excerpts with per-sample metrics (PPL and entropy) for the 1024-step teacher and DiDi-Instruct students from 8 to 128 NFEs, which confirms a clear improvement in narrative quality. The 8-step student exhibits expected repetition, a common artifact of rapid sampling. However, this issue is resolved by 16 NFEs, at which point the student model already surpasses the 1024-step teacher in paragraph-level coherence and topic adherence. This ability to construct focused and specific narratives strengthens as NFEs increase to 128, indicating that DiDi-Instruct not only preserves fluency, but actively enhances the model’s ability to generate structured and coherent text.

### v3

@@ 3.1 Discrete Diffusion Divergence Instruction
A primary challenge in distilling MDMs is the discrete nature of the state space. The gradient formulation in Luo et al. (2023b) relies on differentiating through $\mathbf{z}_{t}$. In MDMs, the forward process involves non-differentiable operations (e.g., $\operatorname*{arg\,max}$), making this gradient inapplicable (Zekri and Boullé, 2025). To address this, Zhu et al. (2025b) proposed a proxy model to approximate $\mathbf{z}_{t}$ and reported promising results for text-to-image tasks. However, its effectiveness in text generation remains underexplored. Instead, we draw inspiration from the policy gradient method (Schulman et al., 2017; Fan et al., 2023) to obtain a rigorous solution. By decomposing the objective gradient as a score function weighted by a log-density ratio, we derive the following gradient estimation.
@@
Let the objective $\mathcal{L}(\nu)$ be the weighted IKL divergences between student and teacher marginals, $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$. The gradient of the objective admits:
@@
where time $t$ is sampled from distribution $\pi(t)$, and $R(\mathbf{z}_{t},t):=\log\mathbf{q}_{\nu}(\mathbf{z}_{t},t)-\log\mathbf{q}_{\theta}(\mathbf{z}_{t},t)$ is the reward (i.e., the log-density ratio between the student and teacher) evaluated at $\mathbf{z}_{t}$.
The proof is deferred to Appendix B. Theorem 3.1 shows that the IKL gradient admits a score-function form that does not differentiate through the discrete sampling path. We sample $\mathbf{x}\sim\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)$, corrupt it to $\mathbf{z}_{t}\sim\mathcal{Q}$, and differentiate only $\log\mathbf{p}_{\nu}(\mathbf{z}_{t}=\mathbf{m},t=1)$. The reward $R(\mathbf{z}_{t},t)$ weights the score to guide distillation toward matching the teacher’s marginal at $t$.
@@ 3.2 Density‑Ratio Estimation
The reward is a mathematically natural choice for distribution matching. However, both $\log\mathbf{q}_{\nu}$ and $\log\mathbf{q}_{\theta}$ are intractable to compute directly. To address this, we avoid estimating the individual densities and instead approximate their ratio. Inspired by Goodfellow et al. (2014); Wang et al. (2025), we train an auxiliary discriminator for this purpose. Intuitively, the optimal discriminator, which differentiates $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$, implicitly encodes their density ratio, as its output probability directly reflects the relative likelihood of the two distributions. We thus model the reward based on the discriminator.
@@
Let $D_{\lambda}:\mathcal{V}^{L}\times[0,1]\to(0,1)^{L}$ be a parameterized discriminator to distinguish samples from the student marginal $\mathbf{q}_{\nu}$ and the teacher marginal $\mathbf{q}_{\theta}$. For the optimal discriminator $D_{\lambda^{\star}}(\mathbf{z}_{t},t)$, the density ratio satisfies
@@
Following Lemma 3.2, we construct a tractable reward signal $R(\mathbf{z}_{t},t)$ by aggregating the log-density ratios across all masked positions. Let $M$ denote the number of masked tokens in the sequence. The reward can be estimated using the discriminator $D_{\lambda}$ as follows:
@@
where $D_{\lambda}^{\ell}(\cdot,\cdot)$ denotes the $\ell$-th element of $D_{\lambda}(\cdot,\cdot)$. This formulation provides a tractable and computable reward signal for our objective based on $D_{\lambda}$. For more details, please check Appendix C.
@@ 3.3 Grouped Reward Normalization
Although the reward estimator in (6) is tractable, its direct use in score-function gradients can exhibit high variance (Williams, 1992). We therefore adopt Group Relative Policy Optimization (Shao et al., 2024; Team et al., 2025) to standardize rewards within each mini-batch. Given a mini-batch of size $G$, we draw $\{t_{i}\}_{i=1}^{G}$ from $\pi(t)$, estimate $\{\mathbf{z}_{t_{i}}\}_{i=1}^{G}$ and $\{R(\mathbf{z}_{t_{i}},t_{i})\}_{i=1}^{G}$ correspondingly [^3] For clarity and concision, we denote $R(\mathbf{z}_{t_{i}}^{1:L},t_{i})$ as $R_{i}$, and $D_{\lambda}(\mathbf{z}_{t_{i}}^{1:L},t_{i})$ as $D_{i}$ throughout.. The final stabilized reward $\widetilde{R}_{i}$ is obtained by normalizing $R_{i}$ with the mean $\mu_{g}$ and variance $\sigma_{g}^{2}$:
$\widetilde{R}_{i}=\frac{R_{i}-\mu_{g}}{\sigma_{g}+\epsilon},\qquad\text{ where }\mu_{g}=\tfrac{1}{G}\sum_{i=1}^{G}R_{i}\quad\text{ and }\quad\sigma_{g}^{2}=\tfrac{1}{G}\sum_{i=1}^{G}(R_{i}-\mu_{g})^{2}.$ (7)
Here, a small constant $\epsilon>0$ is used for numerical stability. Replacing the raw reward $R_{i}$ with this stabilized reward $\tilde{R}_{i}$ in the gradient estimator yields a more robust and stable update rule.
@@ 3.4 The Proposed Algorithms
Adversarial reward estimation. We train a parameterized discriminator $D_{\lambda}$ to separate corrupted samples from $\mathbf{q}_{\nu}$ and $\mathbf{q}_{\theta}$ at time $t$. For each mini-batch index $i$, we start from the all-masked sequence to sample $\mathbf{x}\sim\mathbf{p}_{\nu}$ and $\mathbf{x}^{\prime}\sim\mathbf{p}_{\theta}$, then corrupt both to the same noise level $t_{i}$ via (2) to obtain $\mathbf{z}_{i}\sim\mathcal{Q}$ and $\mathbf{z}^{\prime}_{i}\sim\,\mathcal{Q}$. The discriminator is optimized with balanced binary cross-entropy:
@@
At optimality (i.e., for $\lambda^{\star}$), the discriminator satisfies $D_{\lambda^{\star}}(\mathbf{z}_{t},t)\approx\frac{\mathbf{q}_{\nu}(\mathbf{z}_{t},t)}{\mathbf{q}_{\nu}(\mathbf{z}_{t},t)+\mathbf{q}_{\theta}(\mathbf{z}_{t},t)}$. This indicates that $\mathrm{logit}\,D_{\lambda^{\star}}$ provides a tractable estimate of the log–density ratio used in (6).
Remark 3.3.
Compared with regression-based distillation (e.g., SDTT and DUO), distillation is more delicate to optimize and requires proper regularization, but it is also crucial for achieving high-fidelity generation with very few diffusion steps (Dhariwal and Nichol, 2021). To promote training stability, we (i) initialize both $\mathbf{p}_{\nu}$ and $D_{\lambda}$ from the pre-trained teacher, (ii) warm up $D_{\lambda}$ while freezing $\mathbf{p}_{\nu}$ for an initial phase, and (iii) clip both rewards and gradients in the policy update to avoid exploding updates. Empirically, we monitor discriminator accuracy and observe that it consistently stays in a non-saturated regime.
@@
Figure 2: The visualized pipeline of DiDi-Instruct (summarized in Algorithm 1). Given a fully masked input $\mathbf{z}_{t}$ ($t=1$), both the student $\mathbf{p}_{\nu}$ and the teacher (assistant model) $\mathbf{p}_{\theta}$ produce clean samples $\mathbf{x}$ and $\mathbf{x}^{\prime}$, which are corrupted at $t_{i}\sim\pi(t)$ to form $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{\prime}$. The discriminator $D_{\lambda}$ is trained to classify these outputs, while its reward signal (6) enables the gradient update (5) for the student. The red line represents the gradient flow for the student’s update step, and the blue line represents the gradient flow for the auxiliary model’s update step.
End-to-end training algorithm. The training procedure of DiDi-Instruct alternates between updating the discriminator $D_{\lambda}$ and the student $\mathbf{p}_{\nu}$, which is summarized in Algorithm 1 and visualized in Figure 2. Each iteration consists of two phases: First, the discriminator is updated to better distinguish between corrupted samples from $\mathbf{p}_{\theta}$ and $\mathbf{p}_{\nu}$; We then apply (5) with the normalized reward and the decomposed score to stabilize gradient updates. This alternating scheme yields a robust few-step student generator that closely matches the teacher’s marginals over different corrupt levels. Algorithm 1 DiDi-Instruct Training for each training step do   Sample $t_{i}\sim\pi(t)$, and    $\mathbf{x}\sim{\color[rgb]{0.70,0.00,0.00}\definecolor[named]{pgfstrokecolor}{rgb}{0.70,0.00,0.00}\mathbf{p}_{\nu}}(\mathbf{z}_{t},t=1)$, $\mathbf{x}^{\prime}\sim{\color[rgb]{0.10,0.20,0.65}\definecolor[named]{pgfstrokecolor}{rgb}{0.10,0.20,0.65}\mathbf{p}_{\theta}}(\mathbf{z}_{t},t=1)$.   Corrupt $\mathbf{x}$, $\mathbf{x}^{\prime}$ to partially masked $\mathbf{z}_{i}$, $\mathbf{z}^{\prime}_{i}$.   Update discriminator $D_{\lambda}$ with (8).   Update student $\mathbf{p}_{\nu}$ with (5)-(7), and (9). return student $\mathbf{p}_{\nu}$ and discriminator $D_{\lambda}$. Algorithm 2 RGAS Inference Initialize $\mathbf{z}_{N}\leftarrow(\mathbf{m},\dots,\mathbf{m})$. for $n=N,\dots,1$ do   if $n$ in early stage then    Sample $\mathbf{z}_{n-1}$ with (10) ($h>0$, $M=1$).   else    Sample $\mathbf{z}_{n-1}$ with (10)-(11) ($h=0$, $M\geq 1$).    Set $\mathbf{x}=\mathbf{z}_{0}$ and return sequence $\mathbf{x}$.
@@
where the logits for tokens to be unmasked are augmented with the reward gradient:
@@
Here, $\mathbf{e}_{z_{n}^{\ell}}\in\mathcal{V}$ denotes the one-hot vector with a $1$ at index $z_{n}^{\ell}$ and 0 elsewhere, and $h$ denotes the tilting scale. RGAS adjusts $h$ and the number of candidates $M$ across the denoising process: For early steps ($t_{n}\approx 1$), we use gradient tilting ($h>0,M=1$) to steer global structure toward high-reward regions. For late steps ($t_{n}\approx 0$), we switch to multi-candidate re-ranking ($h=0,M>1$): we draw $M$ candidates $\{\mathbf{z}_{n-1}^{(m)}\}_{m=1}^{M}$ [^4] For simplicity, we denote the $m$-th candidate of $\mathbf{z}_{t_{n}}^{1:L}$ as $\mathbf{z}_{n}^{(m)}$. from (10) with $h=0$, then select $\mathbf{z}_{n-1}$ according to:
@@
To decompose score estimation in (9) and sample the intermediate state $\mathbf{z}_{i}$ at $t_{i}$, we set $h=0$ and $M=1$ to mitigate the risk of reward hacking (Skalse et al., 2022; Gao et al., 2023). The complete procedure is summarized in Algorithm 2.
@@ 4 Experiments
Experimental setup. We distill a pre-trained teacher model into an efficient few-step student generator with DiDi-Instruct. All models are trained on OWT (Gokaslan et al., 2019). Following standard practices (Sahoo et al., 2025), we tokenize the corpus using the GPT-2 tokenizer, pack sequences to a context length of 1024, and hold out the last 100,000 documents for validation.
The teacher is a 169M parameter MDLM with a Diffusion Transformer (Peebles and Xie, 2023) (12 layers, 12 attention heads, 768 hidden dimension). We pre-trained this model from scratch for 1024 NFEs, achieving a perplexity of 38.53 and an entropy of 5.22. The student model shares an identical architecture to ensure a fair comparison. The reward model is a 131M parameter network based on the same backbone, but with a new randomly initialized classification head. This head consists of two linear layers with spectral normalization and a SiLU activation function. During distillation, the reward model and the student generator are trained adversarially in an alternating fashion.
The distillation process runs 10,000 iterations using the AdamW optimizer with a learning rate of $10^{-6}$ and no warm-up. The teacher model was pre-trained on 8 NVIDIA H100 GPUs. Subsequently, our DiDi-Instruct distillation is highly efficient, requiring only a single H100 GPU. All training procedures leverage bfloat16 for acceleration. Additional details are provided in Appendix D.
@@ 4.1 Experimental Result and Analysis
This section shows that DiDi-Instruct not only achieves state-of-the-art performance in sample quality, particularly in the few-step generation regime, but also offers substantial improvements in training and inference efficiency.
Generation results. We evaluate generative quality by sampling text from the distilled student from $8$ to $128$ NFEs and measuring GPT-2 Large generative PPL and average sequence entropy. Figure 1 shows that DiDi-Instruct consistently outperforms all baselines in PPL across all sampling steps. Notably, with only 16 NFEs, our model’s PPL already surpasses that of the 1024-step teacher model. At 1024 NFEs, DiDi-Instruct achieves a final PPL of 15.62, a reduction of over $24\%$ compared to the strongest baseline. These performance gains are achieved with a negligible loss in diversity; the generative entropy from 8 to 128 NFEs is 5.17, 5.21, 5.18, 5.15, and 5.15, respectively (e.g., 1024-step teacher model is 5.22), indicating sample diversity is well-preserved.
We further assess sample fidelity and diversity using MAUVE (Pillutla et al., 2021) and Self-BLEU (Papineni et al., 2002; Montahaei et al., 2019). In appendix E.1, DiDi-Instruct achieves superior distribution matching (higher MAUVE) and reduced mode collapse (lower Self-BLEU) compared to SDTT and DUO, confirming the improvement comes without compromising generative diversity.
Efficiency-performance tradeoff. DiDi-Instruct offers substantial computational advantages in both training and inference. Our single-round distillation framework completes training in around one H100 GPU hour, in contrast to the 2020+ GPU hours required by multi-round methods (Sahoo et al., 2024; Deschenaux and Gulcehre, 2025). During inference, DiDi-Instruct demonstrates a superior throughput-latency profile. Under a standardized benchmark on a single H100 GPU, it achieves $2366$ tokens/sec, representing a $\mathbf{13.2\times}$ speedup over an AR model of the same size at matched perplexity. We provide detailed benchmarks and setup configurations in Appendix E.2.
@@
Sample quality. Appendix E.10 presents full text excerpts with per-sample metrics (PPL and entropy) for the 1024-step teacher and DiDi-Instruct students from 8 to 128 NFEs, which confirms a clear improvement in narrative quality. The 8-step student exhibits expected repetition, a common artifact of rapid sampling. However, this issue is resolved by 16 NFEs, at which point the student model already surpasses the 1024-step teacher in paragraph-level coherence and topic adherence. This ability to construct focused and specific narratives strengthens as NFEs increase to 128, indicating that DiDi-Instruct not only preserves fluency but actively enhances the model’s ability to generate structured, coherent text.